# Ultraviolet photon production rates of the first stars: Impact on the He II 1640 Ă emission line from primordial star clusters and the 21-cm signal from cosmic dawn

### v1

@@ 5 Impact on the 21 cm signal from cosmic dawn
The impact of Pop III stars on the cosmic 21-cm signal is multifold. During the epoch of cosmic dawn, the Lyman-band continuum flux of Pop III stars plays an important role through the WouthuysenâField (WF) effect, wherein Lyman- photons scatter off neutral hydrogen in the intergalactic medium and act to couple the 21-cm spin temperature to the kinetic temperature of the gas. The scattering of Lyman-line photons also acts to raise the kinetic temperature of the intergalactic medium. At the same time, LW photons dissociate H2 molecules, suppressing further Pop III star formation in molecular cooling dark matter minihaloes. During the epoch of reionization, Pop III stars that are still able to form ionize their surrounding gas, leading to a suppression in the 21-cm signal.
@@
In Fig. 6, we plot the lifetime-integrated Lyman-band photon production rates per ZAMS stellar baryon of our Pop III models in relation to the corresponding predictions presented by Gessey-Jones et al. (2022) for non-rotating Pop III stars, and those of Liu et al. (2025a) for rotating Pop III stars. Both of these are, like the Muspelheim ones, based on TLUSTY stellar atmosphere SEDs, but employ different assumptions concerning stellar evolution. The Gessey-Jones et al. (2022) and Liu et al. (2025a) models both trace evolution up to the end of main sequence under the assumption of no mass loss, whereas our models also include later evolutionary stages and â in the case of Yoon et al. (2012) models with rotation â substantial mass loss. Whereas the Gessey-Jones et al. (2022) models are based on non-rotating stars, the Liu et al. (2025a) models are based on Sibony et al. (2022) tracks for polytrope stars subject to chemically homogeneous evolution due to strong internal mixing. While Sibony et al. (2022) choose to remain agnostic about the physical reason for what would cause this mixing, fast rotation represents a strong candidate, as evident by the rotating Yoon et al. (2012) models, which in many cases also undergo chemically homogeneous evolution. For simplicity, we hereafter refer to the Liu et al. (2025a) models as rotating, in line with the discussion of these models by Liu et al. (2025a, b).
@@
Differences in between the models at any given mass are within about a factor of two, with the rotating stars below 30 Mâ consistently resulting in higher rates compared with their non-rotating counterparts due to their increased lifetimes. Interestingly for our discussion, we note a large difference compared to the Gessey-Jones et al. (2022) models at a mass of 20 Mâ, where the Yoon et al. (2012) models for rotating stars (which have the potential to boost the HeII1640 emission line equivalent width; Fig. 5) give rise to an that is higher by a factor of . Since it is important to establish whether differences of this amplitude could have any significant impact on 21-cm signatures from the pre-reionization Universe, we use the semi-numerical code 21cmSPACE (e.g. Fialkov et al., 2014; Reis et al., 2021; Gessey-Jones et al., 2022, 2025)222https://www.cosmicdawnlab.com/21cmSPACE to simulate the global 21-cm signal and the spherically averaged 21-cm power-spectrum at â40, under the extreme assumption of a single-star initial mass function where all Pop III stars have the same mass (a rough approximation of the case where the initial mass function is strongly peaked at this mass).
@@
The simulations are performed on cosmological box size of , and include effects such as baryon dark-matter streaming velocity, Lyman- heating and scattering and LymanâWerner feedback (important during the cosmic dawn) as well as photoheating feedback, X-ray heating and ionization from flexible spectra models (important during the epoch of reionization). In particular, for this work, we use the code version described in Liu et al. (2025a) where the effect of ionizing photons from Pop III stars was introduced. To put the Yoon et al. (2012) Pop III models into context, we also run 21cmSPACE with the Pop III models of Gessey-Jones et al. (2022) and Liu et al. (2025a).
We consider two sets of astrophysical parameters for comparison:
| Scenario 1: | |||
| Scenario 2: |
@@
where and represent the star formation efficiency for Pop II and III stars, respectively, and is the recovery time for star formation after the onset of Pop III supernovae. Due to the higher ratio and the longer of scenario 2, the impact of the detailed properties of Pop III stars on 21 cm observables during cosmic dawn will be higher for this model compared to that of scenario 1. To pinpoint the role of Lyman-band emissivities of Pop III stars in shaping the 21-cm signal, we switch off X-ray heating from Pop III stars. For Pop II, we adopt an X-ray luminosity tied to the star formation rate (SFR) of each halo, according to erg s-1 M yr-1 (Fragos et al., 2013a, b). Throughout, we assume instantaneous emission from Pop III stars, which is a reasonable assumption for stellar masses (Gessey-Jones et al., 2022).
@@
Fig. 7 shows the resulting global 21-cm signal, and the power spectrum generated at a scale of cMpc-1, for the Yoon et al. (2012) 20 Mâ model with rotation, in comparison with the 20 Mâ model from Gessey-Jones et al. (2022) without rotation, the 20 Mâ from Liu et al. (2025a) with rotation, and the 20 and 200 Mâ Yoon et al. (2012) model without rotation. The 200 Mâ model is included here as an example of a type of star that, because of its high , could produce strong HeII1640 emission despite being non-rotating (Fig. 5).
@@
In the case of the 21-cm power spectra, we also plot the expected 5 detection limit for 1080 hours of tracking-scan observations using the Square Kilometre Array (SKA-Low) with 6 hours of observation per night for 180 days. The derived detection limit considers both the thermal noise and the sample variance. To estimate the uncertainty in the power spectrum due to thermal noise, we used the package 21cmSense333https://github.com/rasg-affiliates/21cmSense/tree/main (Pober et al., 2013, 2014). We only considered the SKA-Low AAâ array configuration (Sridhar et al., 2024) within a radius of 2 km from the array center and assumed the moderate foreground model in 21cmSense. The sample variance is calculated according to the equation
@@ 5.1 The global 21-cm signal
The differences in the global signal between the Pop III stellar models can be explained by considering the epochs where the early onset of the global signal absorption trough is almost entirely dictated by the Lyman band emissivities (Fig. 6)444Although LymanâWerner photons do contribute to suppression of star-formation in molecular cooling haloes, and we account for this., while the absorption trough at is also sensitive to the ionizing flux from the stars (Fig. 3).
During the cosmic dawn (), a higher Lyman-band flux leads to a stronger WF effect and more rapid onset of significant 21-cm absorption. This would be strongest for the rotating 20 Mâ Yoon et al. (2012) model and weakest for the non-rotating 20 Mâ Gessey-Jones et al. (2022) and non-rotating 200 Mâ Yoon et al. (2012) model. Since the latter two exhibit similar Lyman fluxes, they have nearly identical global signal absorption amplitude at in both scenarios 1 and 2. The rotating 20 Mâ Liu et al. (2025a) model and non-rotating 20 MâYoon et al. (2012) model fall in between these extremes and approximately similar behaviour. The difference in the 21-cm signal at these high redshifts () is of the order of few mK in scenario 1 and up to at its largest in scenario 2.
During the epoch of heating and reionization (), the differences in these models is dominated by the ionization from Pop III stars (as explored in Liu et al., 2025a). In this case, the rotating 20 Mâ stellar models from Yoon et al. (2012) and Liu et al. (2025a) predict similarly strong ionizing fluxes which manifest as weak emission signal at for scenario 1, and a suppressed absorption trough at for scenario 2. Indeed, with these more extreme stellar models, there is already a 10â20 per cent reduction in the neutral hydrogen fraction at (see Fig. 8), and a proportional suppression of the signal. This is also true for the case of the non-rotating 200 Mâ Yoon et al. (2012) model, while the 20 Mâ non-rotating models of Yoon et al. (2012) and Gessey-Jones et al. (2022) produce deeper troughs due to late reionization.
@@ 5.2 The 21-cm power spectrum
The predicted high-redshift 21-cm power spectra at , for both scenarios 1 and 2, fall below the detection limit of the SKA. At lower redshifts, where the power spectrum is deemed detectable according to our error model, the differences between models are marginal in scenario 1. Still, they are more pronounced and potentially detectable in scenario 2. The largest difference comes from the ionization signature of rotating Pop III stars. The 20 Mâ Yoon et al. (2012) and Liu et al. (2025a) models, which both undergo chemically homogenous evolution, produce power spectra which deviate from those predicted for the non-rotating 20 Mâ models of Gessey-Jones et al. (2022) and Yoon et al. (2012). In both cases, the emission signal vanishes as the Universe is rapidly ionized, and the ionized bubbles imprint detectable fluctuations in the power spectrum at . Here too, the non-rotating 200 Mâ Yoon et al. (2012) model shows an intermediate behaviour between the two cases.
Overall, we conclude that the dramatic difference seen between the He+-ionizing rates (Fig. 4) and HeII1640 equivalent widths (Fig. 5) of rotating vs. non-rotating stars do not necessarily translate into equally dramatic 21-cm signatures. While differences due to rotation may be imprinted in the 21-cm power spectrum detectable by SKA, these are only conspicuous in some parts of the Pop III parameter space (i.e. in our scenario 2 where it boosts the ionization peak by factor of a 2â3, but not in scenario 1).
@@ 6.1 Blackbody versus stellar atmosphere SEDs
When comparing our results with the models presented by Klessen & Glover (2023), based on blackbody spectra, we find in Fig. 3 the largest discrepancy in He+ ionizing rates for stars , where blackbody spectra result in rates orders of magnitude higher. These differences, also echoed in our Murphy et al. (2021a) comparison of momentary for stellar atmosphere spectra versus blackbody spectra in Fig. 2, are due to the different shapes of blackbody and stellar atmosphere SEDs and the lack of continuum breaks in the former case, as shown in Fig. 1. Hence, blackbody spectra should not be used to estimate Pop III photon production rates or fluxes at such high energies.
@@
We also note a significant deviation in all UV bands for stars when comparing blackbody and stellar atmospheres, as shown in Fig. 3. While Murphy et al. (2021b) did comment on the reliability of the blackbody approximation for the Murphy et al. (2021b) tracks (through a comparison to the results of Schaerer 2002), this offset was not noted since the Murphy et al. (2021b) study of UV photon production rates considered only stars of masses from the Murphy et al. (2021a) set. The blackbody approximation from their lower-mass tracks is adopted by Klessen & Glover (2023) which overestimate the ionizing rates for all low mass stars and slightly underestimate the LW rates for these same stars. However, as long as the Pop III IMF is assumed to be top-heavy, this discrepancy should not be a major source of error, as the ionizing photon production in that case would be dominated by the more massive stars.
6.2 The LymanâWerner emissivities of Pop III stars
@@
The LW photon production rates of Pop III stars are important for establishing the LW radiation background during cosmic dawn and for the quenching of subsequent star formation in low-mass haloes. When the lifetime-integrated is expressed as LW photons per stellar baryon, there are two distinct values that are frequently featured in the literature: 4800 and . Our models do not favour either of these, but instead favour values in the intermediate range to (see Fig. 3; values in this figure can be converted to log LW photons per stellar baryon by subtracting ). The lower literature value (4800) stems from Barkana & Loeb (2005), based on Pop III spectra from Bromm et al. (2001) and an adopted power-law slope of the spectrum, whereas the higher value () is usually attributed to Schaerer (2002). Indeed, when the Schaerer (2002) Table 4 values are converted into these units, they peak at LW photons per baryon at a Pop III stellar mass of 60 Mâ. However, as shown in Fig.  3, the Schaerer (2002) are higher than ours at by up to a factor (0.7 dex). Mas-Ribas et al. (2016) also note difficulties in reproducing the Schaerer (2002) LW rates and suggest that the Schaerer (2002) LW rates may have been integrated from 11.2 eV (wavelength 1107 Ă ) to infinity, rather than to 13.6 eV (wavelength 912 Ă ), which would lead to boosted rates. Lifetime-integrated LW rates per stellar baryons in the range for Pop III stars are also reported in the recent paper by Costa et al. (2025), and in that case the LW rates were indeed integrated from 11.2 eV to infinity (Guglielmo Costa, private communication). After correcting the Costa et al. (2025) limit from 11.2 to 13.6 eV, their values agree with ours. Our values are moreover in good agreement with those recently presented by Liu et al. (2025a) for the non-rotating Pop III tracks of Gessey-Jones et al. (2022) and the rotating Pop III tracks of Sibony et al. (2022). Hence, our recommendation is to use values for the lifetime-integrated LW photon counts per stellar baryons of to .

### v2

@@ 5 Impact on the 21 cm signal from cosmic dawn
The impact of Pop III stars on the cosmic 21-cm signal is multifold. During the epoch of cosmic dawn, the Lyman-band continuum flux of Pop III stars plays an important role through the Wouthuysen–Field (WF) effect, wherein Lyman- photons scatter off neutral hydrogen in the intergalactic medium and act to couple the 21-cm spin temperature to the kinetic temperature of the gas (Wouthuysen, 1952; Field, 1958). The scattering of Lyman-line photons also acts to raise the kinetic temperature of the intergalactic medium (e.g. Chuzhoy and Shapiro, 2007; Reis et al., 2021). At the same time, LW photons dissociate H2 molecules, suppressing further Pop III star formation in molecular cooling dark matter minihaloes (e.g. Fialkov et al., 2013; Visbal et al., 2014; Mirocha et al., 2018). During the epoch of reionization, Pop III stars that are still able to form ionize their surrounding gas, leading to a suppression in the 21-cm signal (e.g. Liu et al., 2025a)
@@
In Fig. 6, we plot the lifetime-integrated Lyman-band photon production rates per ZAMS stellar baryon of our Pop III models in relation to the corresponding predictions presented by Gessey-Jones et al. (2022) for non-rotating Pop III stars, and those of Liu et al. (2025a) for rotating Pop III stars. Both of these are, like the Muspelheim ones, based on TLUSTY stellar atmosphere SEDs, but employ different assumptions concerning stellar evolution. The Gessey-Jones et al. (2022) and Liu et al. (2025a) models both trace evolution up to the end of main sequence under the assumption of no mass loss, whereas our models also include later evolutionary stages and – in the case of Yoon et al. (2012) models with rotation – substantial mass loss. Whereas the Gessey-Jones et al. (2022) models are based on non-rotating stars, the Liu et al. (2025a) models are based on Sibony et al. (2022) tracks for polytrope stars subject to chemically homogeneous evolution due to strong internal mixing. While Sibony et al. (2022) choose to remain agnostic about the physical reason for what would cause this mixing, fast rotation represents a strong candidate, as evident by the rotating Yoon et al. (2012) models, which in many cases also undergo chemically homogeneous evolution. For simplicity, we hereafter refer to the Liu et al. (2025a) models as rotating, in line with the discussion of these models by Liu et al. (2025a, b).
@@
Differences in between the models at any given mass are within about a factor of two, with the rotating stars below 30 M⊙ consistently resulting in higher rates compared with their non-rotating counterparts due to their increased lifetimes. Interestingly for our discussion, we note a large difference compared to the Gessey-Jones et al. (2022) models at a mass of 20 M⊙, where the Yoon et al. (2012) models for rotating stars (which have the potential to boost the HeII1640 emission line equivalent width; Fig. 5) give rise to an that is higher by a factor of . Since it is important to establish whether differences of this amplitude could have any significant impact on 21-cm signatures from the pre-reionization Universe, we use the semi-numerical code 21cmSPACE (e.g. Fialkov et al., 2014; Reis et al., 2021; Gessey-Jones et al., 2022, 2025)222https://www.cosmicdawnlab.com/21cmSPACE to simulate the global 21-cm signal and the spherically averaged 21-cm power-spectrum at , under the extreme assumption of a single-star IMF where all Pop III stars have the same mass (a rough approximation of the case where the IMF is strongly peaked at this mass).
@@
The simulations are performed on a cosmological box size of , and include effects such as baryon dark-matter streaming velocity (Fialkov et al., 2012), Lyman–Werner feedback (important during the cosmic dawn; Fialkov et al., 2013; Muñoz et al., 2022), Lyman- heating and scattering (Reis et al., 2021), photoheating feedback (Sobacchi and Mesinger, 2013; Cohen et al., 2016), X-ray heating (Fialkov et al., 2014; Fialkov and Barkana, 2014) and ionization from flexible spectra models (important during the epoch of reionization). In particular, for this work, we use the code version described in Liu et al. (2025a) where the effect of ionizing photons from Pop III stars was introduced. To put the Yoon et al. (2012) Pop III models into context, we also run 21cmSPACE with the Pop III models of Gessey-Jones et al. (2022) and Liu et al. (2025a).
@@
We consider two models, with varying Pop III star formation efficiencies (i.e. fraction of gas converted in stars): Model 1 where Pop II SFR dominates at with and Model 2 where Pop III SFR is comparable to Pop II at with , as shown in Figure 8. In both models, we fix the Pop II star-formation efficiency to which generally matches observations at UV luminosity function observations at (Tacchella et al., 2018; Sipple and Lidz, 2024; Dhandha et al., 2025), and an star-formation averaging timescale of . We also fix Myr, which is the recovery time for Pop II star formation after the onset of Pop III supernovae (Magg et al., 2022a). To pinpoint the role of Lyman-band emissivities of Pop III stars in shaping the 21-cm signal, we switch off X-ray heating from Pop III stars. For Pop II, we adopt an X-ray luminosity tied to the star formation rate (SFR) of each halo, according to erg s-1 M yr-1 (Fragos et al., 2013b, a), which is consistent with constraints on the diffuse X-ray background, and HERA constraints on the 21-cm power spectrum at (Dhandha et al., 2025). Throughout, we assume instantaneous emission from Pop III stars, which is a reasonable assumption for stellar masses (Gessey-Jones et al., 2022).
@@
Fig. 7 shows the resulting global 21-cm signal, and the power spectrum generated at a scale of cMpc-1, for six models: four models assuming a single-mass IMF of 20 M⊙ for the rotating and non-rotating Yoon et al. (2012) model, non-rotating Gessey-Jones et al. (2022) model, and a rotating Liu et al. (2025a) model, as well as two additional models of single-mass IMFs at 200 M⊙ for rotating and non-rotating Yoon et al. (2012) stars. The 200 M⊙ non-rotating Yoon et al. (2012) model is included here as an example of a type of star that, because of its high , could produce strong HeII1640 emission despite being non-rotating (Fig. 5).
@@
In the case of the 21-cm power spectra, we also plot the expected 5 detection limit for 1080 hours of tracking-scan observations using the Square Kilometre Array (SKA-Low) with 6 hours of observation per night for 180 days. The derived detection limit considers both the thermal noise and the sample variance. To estimate the uncertainty in the power spectrum due to thermal noise, we used the package 21cmSense333https://github.com/rasg-affiliates/21cmSense/tree/main (Pober et al., 2013, 2014). We only considered the SKA-Low AA∗ array configuration (Sridhar et al., 2024) within a radius of 2 km from the array center and assumed the moderate foreground model in 21cmSense. The sample variance is calculated according to the equation
@@ 5.1 The global 21-cm signal
The differences in the global signal between the Pop III stellar models can be explained by considering the epochs where the early onset of the global signal absorption trough is almost entirely dictated by the Lyman band emissivities (Fig. 6)444Although Lyman–Werner photons do contribute to suppression of star-formation in molecular cooling haloes, and we account for this., while the absorption trough at is also sensitive to the ionizing flux from the stars (Fig. 3).
During the cosmic dawn (), a higher Lyman-band flux leads to a stronger WF effect and more rapid onset of significant 21-cm absorption. This would be strongest for the rotating 20 M⊙ Yoon et al. (2012) model and weakest for the rotating 200 M⊙ Yoon et al. (2012) model, which indicates that the 21-cm signal observations can have the potential to distinguish the IMF of stellar populations (e.g. as also demonstrated in Gessey-Jones et al., 2025). The signal also shows differentiating features between the rotating and non-rotating 200 M⊙ Yoon et al. (2012) models, and a similar (but reversed) trend for the 20 M⊙ Yoon et al. (2012) models. In case of the rotating 20 M⊙ Liu et al. (2025a) model, since it exhibits a similar Lyman flux to the non-rotating 20 M⊙ Yoon et al. (2012), it shows similar behaviour at high redshifts. The difference in the 21-cm signal is of order few mK at the onset of the absorption signal up to () when Lyman-coupling is efficient (at for Model 1 and for Model 2).
During the epoch of heating and reionization (), the differences in these models is dominated by the ionization from Pop III stars (as explored in Liu et al., 2025a). The case of rotating models show convergence since they predict similarly strong ionizing fluxes across stellar masses (at M⊙), manifested as a suppressed absorption through in both Model 1 and Model 2. In the latter case, the differences are particularly large with a difference also seen in the redshift of the absorption signal as ionization becomes efficient. Indeed, this is reflected in the neutral hydrogen fraction shown in Fig. 9 with a reduction between the non-rotating and rotating cases (and thus, a proportional suppression of the signal). With the non-rotating models, the 20 M⊙ Gessey-Jones et al. (2022) model has the least efficient ionization, followed by 20 M⊙ Yoon et al. (2012) model (which shows similar absorption amplitude/timing), and lastly the 200. M⊙ Yoon et al. (2012) model.
@@ 5.2 The 21-cm power spectrum
The predicted high-redshift () 21-cm power spectra for Model 1 falls below the detection limit of the SKA, and close to the limit for Model 2. At lower redshifts (), where the power spectrum is deemed detectable according to our error model, the differences between stellar models are marginal in Model 1 (especially if considering degeneracies with other astrophysical parameters that we keep fixed here). However, there are more pronounced and detectable differences in Model 2 which come from the ionization signature of the rotating Pop III stars, as in the global signal. In these cases, as the Universe is rapidly ionized, the ionized bubbles imprint boosted fluctuations in the power spectrum at . The non-rotating cases show a similar trend to the global signals, where the 20 M⊙ Gessey-Jones et al. (2022) model shows the smallest ionization peak of mK2, while the 20 M⊙ Yoon et al. (2012) model and 200 M⊙ Yoon et al. (2012) model show higher peaks by and .
Overall, we conclude that the dramatic difference seen between the He+-ionizing rates (Fig. 4) and HeII1640 equivalent widths (Fig. 5) can imprint some detectable 21-cm signatures. The differences are modest for low Pop III star-formation efficiency (Model 1), potentially only detectable with high-accuracy global signal experiments (noise levels of mK) at . For high Pop III star-formation efficiency (Model 2, which shows early reionization), the differences imprinted in the 21-cm power spectrum at are more dramatic and detectable by SKA.
@@ 6.1 Blackbody versus stellar atmosphere SEDs
When comparing our results with the models based on blackbody spectra presented by Klessen and Glover (2023) in Fig. 3, we find a large discrepancy is in the He+ ionizing rates for stars , where blackbody spectra result in rates orders of magnitude higher than those based on stellar atmosphere spectra. These differences, also echoed in our Murphy et al. (2021a) comparison of momentary for stellar atmosphere spectra versus blackbody spectra in Fig. 2, are due to the different shapes of blackbody and stellar atmosphere SEDs and the lack of continuum breaks in the former case, as shown in Fig. 1. Hence, blackbody spectra should not be used to estimate Pop III photon production rates or fluxes in the He+ ionizing energy range.
@@
We also note significant deviations in the other UV bands for stars . While Murphy et al. (2021b) did comment on the reliability of the blackbody approximation for the Murphy et al. (2021b) tracks (through a comparison to the results of Schaerer 2002), this offset was not noted since the Murphy et al. (2021b) study of UV photon production rates considered only stars of masses from the Murphy et al. (2021a) set. The blackbody approximation from their lower-mass tracks is adopted by Klessen and Glover (2023) which overestimates the ionizing rates for all low mass stars and slightly underestimates the LW rates for these same stars. However, as long as the Pop III IMF has a characteristic mass – which is often argued to be the case (e.g. Hartwig et al., 2024; Liu et al., 2024) – this discrepancy should not be a major source of error, as the ionizing photon production in that case would be dominated by the more massive stars.
6.2 The Lyman–Werner emissivities of Pop III stars
@@
The LW photon production rates of Pop III stars are important for establishing the LW radiation background during cosmic dawn and for the quenching of subsequent star formation in low-mass haloes. When the lifetime-integrated is expressed as LW photons per stellar baryon, there are two distinct values that are frequently featured in the literature: 4800 and . Our models do not favour either of these, but instead favour values in the intermediate range to (see Fig. 3; values in this figure can be converted to log LW photons per stellar baryon by subtracting ). The lower literature value (4800) stems from Barkana and Loeb (2005), based on Pop III spectra from Bromm et al. (2001) and an adopted power-law slope of the spectrum, whereas the higher value () is usually attributed to Schaerer (2002). Indeed, when the Schaerer (2002) Table 4 values are converted into these units, they peak at LW photons per baryon at a Pop III stellar mass of 60 M⊙. However, as shown in Fig. 3, the Schaerer (2002) are higher than ours at by up to a factor (0.7 dex). Mas-Ribas et al. (2016) also note difficulties in reproducing the Schaerer (2002) LW rates and suggest that the Schaerer (2002) LW rates may have been integrated from 11.2 eV (wavelength 1107 Å) to infinity, rather than to 13.6 eV (wavelength 912 Å), which would lead to boosted rates. Lifetime-integrated LW rates per stellar baryons in the range for Pop III stars are also reported in the recent paper by Costa et al. (2025), and in that case the LW rates were indeed integrated from 11.2 eV to infinity (Guglielmo Costa, private communication). After correcting the Costa et al. (2025) limit from 11.2 to 13.6 eV, their values agree with ours. Our values are moreover in good agreement with those recently presented by Liu et al. (2025a) for the non-rotating Pop III tracks of Gessey-Jones et al. (2022) and the rotating Pop III tracks of Sibony et al. (2022). Hence, our recommendation is to use values for the lifetime-integrated LW photon counts per stellar baryons of to .